solve each system of equations.Let Z be the parameter.
x + 2y +3z =11
2x - y + z = 2

Answer by stanbon(75887)

solve each system of equations.Let Z be the parameter.
x + 2y +3z =11
2x - y + z = 2
---------------
x + 2y = -3z+11
2x - y = -z+2
----------------
Let A be the coefficient matrix of x and y
Then A^-1 = [.2 .4]
[.4 -.2]
--------
Let B = [-3z+11]
[-z+2 ]
-------------------
Then [A^-1][B]= (1/5) [-3z+11-2z+4]
[-6z+22+z-2 ]
=[-z+3]
[-z+4]
===============
So x= -z+3 and y=-z+4
=================
Cheers,
Stan H.
------------------------